OpenStudy (anonymous):

If you perform 10 joules of work lifting a 10-N box from the floor to a shelf, how high is the shelf? A. 1 m B. 2 m C. 10 m D. 100 m

hartnn (hartnn):

Joule = newton * Meter 10 = 10* (???)

OpenStudy (anonymous):

I have no clue. I don't understand how to even begin doing this problem.

OpenStudy (anonymous):

@jim_thompson5910 Any way you could help me here?

jimthompson5910 (jim_thompson5910):

Work = Force * Displacement

jimthompson5910 (jim_thompson5910):

W = F*d 10 = 10*d solve for d

OpenStudy (anonymous):

So D is 1?

OpenStudy (unklerhaukus):

The work done \(W\) in joules against gravity, equals the weight force \(F\)in newtons time the vertical displacement \(y\) in meters (lifting something up takes positive work) \[W=Fy\]

jimthompson5910 (jim_thompson5910):

correct, displacement is 1 meter

jimthompson5910 (jim_thompson5910):

since d = 1

OpenStudy (anonymous):

Okay. Thank you so much. (:

jimthompson5910 (jim_thompson5910):

yw